1) Your company needs to borrow funds and has several options available to it, Loans A, B and C. The interest rates (APR) for these options are given below. What is the EAR of the loan option the company should choose?

Loan A (APR, compounding frequency) 5.95%, semi-annually // Loan B (APR, compounding frequency) 6.02%, monthly // Loan C (APR, compounding frequency) 5.95%, quarterly

2) Your company has an issue of $100 par value annual coupon bonds with 7 years remaining until maturity. The annual coupon rate is 3.45%, with a $96.60 current price of the bonds . What is the yield to maturity on the bonds?

3) Your company has an issue of $1,000 par value bonds that offer an 8% coupon rate paid semi-annually. The bonds have 5 years remaining until maturity. The market’s required return on these bonds is 3.15%. What is the amount of each coupon payment?
Answer #1

1) EAR of Loan A =(1+APR/2)^2-1 =(1+5.95%/2)^2-1 =6.04%
EAR of Loan B =(1+APR/12)^12-1 =(1+6.02%/12)^12-1 =6.19%
EAR of Loan C =(1+APR/4)^4-1 =(1+5.95%/4)^4-1 =6.08%

2) PV of Bond =96.60
Par Value =100
Number of Years =7
Coupon =Coupon Rate*Par Value =3.45%*100 =3.45
YTM using Financial Calculator
N=7;PMT=3.45;PV=-96.60;FV=100;CPT I/Y =4.02%

3)
Amount for each Coupon =Coupon rate*Par Value/2 =8%*1000/2 =40